# Position paper: Do not explain (vision models) without context

### v2

@@ 5 XAI failures
We provide examples when popular XAI techniques fail to explain in a comprehensive way the predictions of the model when spatial context is crucial. We fine-tuned Resnet-50 and ViT base models in low data regime as it is done in work (Jia et al., 2022) using Visual Task Adaptation Benchmark (VTAB) (Zhai et al., 2019). The models were pretrained on Imagenet (Resnet - in a supervised manner), ViT - in a contrastive manner (Moco v3). We focused on the subset of VTAB called ‘structured’ where the labels depend on spatial context. In the experiments, we use two datasets: KITTI (Geiger et al., 2012) where images were collected using sensors in the car - the task is to predict the binned distance to the closest vehicle in the scene, dsprites (Matthey et al., 2017) where images of simple shapes undergo rotations and other shifts in the space - the task is to predict binned orientation. Hence, we analyse the only real-life dataset in VTAB and one of the few synthetic datasets.
Having fine-tuned models of a satisfactory performance (similar to the one claimed in (Jia et al., 2022)), we applied 5 different XAI techniques: GradientSHAP (Lundberg & Lee, 2017), IntegratedGradients (Sundararajan et al., 2017), Occlusion (Zeiler & Fergus, 2014), Saliency (Simonyan et al., 2013), LIME (Ribeiro et al., 2016) to check if they managed to explain the model’s decisions correctly. In Figure 6, we provide images with the models’ explanations. The colours reflect the extent to which particular parts contribute to the final models’ decisions (green - means positive attribution, whereas, red - negative). We can see that the existing XAI methods fail to explain the correct model decisions i.e. that the nearest vehicle is within the distance of 8 to 20 meters.
The explanations fail - they are inaccurate, vague and difficult to interpret. It can be seen that the nearest vehicle is highlighted in green (using the LIME method), however, it is not clear that the distance is crucial. One may conclude that the same explanations could be seen also in the object recognition tasks i.e. cars vs. planes. Therefore, it may be important to design spatial XAI techniques that would be tailored to different ML tasks or context types.
@@ 6 Spatial context in XAI
The topic of spatial context was considered at the stage of a training of neural networks but it was of much smaller interest in XAI. We showed that there is still plenty to do as popular XAI methods fail. Note that without proper XAI tools, we would not be able to explain the predictions of CNNs and CapsuleNets when the same features are within input images but placed differently in the scene. The existing methods in the case of ‘deformed face’ (Figure 5) will most probably highlight the same key elements of the image both in the case of CNNs and CapsuleNets even though the two DL models will return different predictions.
The investigation of whether the spatial context plays a role and to what extent in the decision-making of DL models is a niche that should be fulfilled. We call this area of research spatial XAI. There are only a few works in this domain. As a result of the analyses, we distinguish four main approaches: intrinsically explainable models, methods introducing measures of spatial context, leveraging current XAI methods and the ones based on input-output relationships.
@@
The work (Bardool et al., 2019) encompasses two aspects as it proposes a new explainable-by-design architecture for detection and the metrics to evaluate the importance of context. The authors modify the well-known detector architecture adding a contextual learner to extract contextual cues from the scenes. The contextual feature maps for each class give insights into the way the model benefits from context. The other contribution is the two scoring functions to measure the contextual relevance of detections in relation to other objects present in the image and the general scene. The first score is calculated as a summation of the elements of the feature map corresponding to a bounding box within the mask region, divided by the area of the bounding box’s mask. The second scoring method is analogous but it is calculated based on contextual feature maps generated for each class instead of bounding boxes. The metrics are later used for the comparison of different DL models. The authors challenged the proposed DL model on a set of hypotheses referring to spatial context. It was possible thanks to the fact that they had control over the generation of a synthetic dataset so that you could prepare samples violating the underlying spatial rules in the dataset and see how the model behaves in such scenarios.
@@
Sacha et al. (2023) analyse the spatial aspect within explainable-by-design prototypical models. They define the problem of ‘spatial explanation misalignment’ and introduce the metric to quantify this phenomenon. Moreover, they propose a benchmark and a method to mitigate this spatial misalignment in this particular kind of models. It was observed that sometimes the location of the explanation within the input image changes when the input image is modified in regions that should not influence the model prediction (such image modifications are not in meaningful areas and can be thought of as distractors).
Tomaszewska et al. (2024) focuses on metrics to evaluate the preservation of spatial context within DL models. The authors proposed three spatial context measures (, , ). The quantitative measurement is possible thanks to the incorporation of spatial regression models which are run on patch features to predict the respective attention scores assigned by the DL models. The method is general and can be applied to any attention-based vision model that most often finds application in the analysis of high-resolution images. In the work, the method is illustrated in the histopathology use case where the spatial context in the input is understood as spatial autocorrelation between patches of Whole Slide Images.
@@
One of the early XAI methods is LIME (Ribeiro et al., 2016) which operates on superpixels from input images assigning them the importance score explaining the model’s decision. In work, (Rabold et al., 2020), the authors propose to build upon the LIME method by incorporating logic rules obtained by the Inductive Logic Programming system, Aleph. The authors, therefore, propose to combine visual and symbolic explanations to explain spatial relationships.
@@
In Deep Learning, often ablation studies are performed where the particular parts of the networks are switched off to check whether they are necessary and what their role is. This evaluation is often performed by comparing the performance of the original network and the modified one. In XAI, a similar technique is applied to images that undergo some modifications. This is done in so-called permutation-based XAI methods where parts of the input image are erased to see how it impacts predictions. In work, Tomaszewska et al. (2023), an analogous approach was used to evaluate the importance of neighborhood. The model for prediction of whether the patch of histopathology contains tumor or not was trained on images containing the region of interest and neighborhood. The ground truth label is provided only based on the region of interest. Later, the inference was performed with images with a gradually decreased size of the neighborhood. It was analysed how much the initial performance metrics were impacted by the limitation of contextual information. In the case of a so-defined neighborbood, a simple application of the existing XAI techniques could show whether models pay attention to the neighborhood or not.
Another input-output relationship approach accompanied with a new XAI method is proposed by Haldekar et al., 2017. The authors cover some parts of the input image (using a sliding window) and later compute the influence it has on a value of cross entropy compared to the situation when no modifications to the input image are made. As a result, the heatmaps showing key parts of the images are generated. The authors present some heatmaps visualizing the relationships (below, behind and above). However, the limitation of the provided heatmaps is the fact that without the authors’ captions, it would be difficult to decipher what type of relationships are highlighted. Moreover, the authors analyse by zeroing out) whether there are specific groups of neurons in MLP layers responsible for detecting particular spatial relationships.
Analysis of spatial relation understanding is performed in text-image generation models where it is assessed if the spatial relations specified in the prompt are preserved within generated images (Cho et al., 2022).
@@ 7 Needed research directions
In some of the mentioned papers on spatial context (Haldekar et al., 2017; Bardool et al., 2019; Swingler & Bath, 2020; Rabold et al., 2020), the claims were investigated on synthetically generated datasets in order to have better control over the experiment setup which facilitates verification of research hypotheses. These datasets were not made publicly available and, moreover, were not of sophisticated compositionality as there were only some simple objects in relative relationships put on the plain background. With the advent of the diffusion–based image generative models i.e. stable diffusion (Rombach et al., 2021), it is possible to generate more realistic images with the desired spatial relationships between objects (samples in Figure 3).
Assessment of spatial reasoning skills in image-generation models
PAINTSKILLS (Cho et al., 2022) is a compositional diagnostic evaluation dataset where images are generated from prompts containing information on spatial relationships (above, below, left, right) between objects. Note that these relationships are relatively simple compared to the ones we outlined in Section 2.2. The dataset is used as one of the scenarios verifying reasoning skills within the recently proposed benchmark, Holistic Evaluation of Text-to-Image Models (HEIM) (Lee et al., 2023). Within the reasoning scenario, it is checked whether models understand objects, counts, and spatial relations (compositionality) within images. Some task-specific benchmarks dedicated to the evaluation of the vision-language model’s ability to understand scene composition were proposed (Ma et al., 2022; Ray et al., 2023).
@@
We believe that analogous benchmarks should be proposed to check the reasoning of other types of vision models with a focus on assessing whether the spatial context is properly taken into account. To the best of our knowledge, there is no such benchmark even though there are many vision models that benefit from contextual information (Section 4). Such models are given an image as an input as opposed to diffusion models that get a prompt and therefore, different methods for their investigation should be applied. The analysis of reasoning capabilities is performed in the case of models solving IQ tests but not the real-world cases that we point out.
@@
Such a benchmark incorporating spatial relations understanding could serve as an important point in a checklist before the launch of DL models in critical domains. Such a criterion could fulfil some of the requirements of ensuring models’ safety. For this benchmark to be successful, some new metrics to evaluate to what extent the models take into account spatial context should be proposed. So far models’ investigation mostly relies on visual XAI methods (heatmaps) which are not a sufficient way to explain models in cases when spatial context matters as shown in Figure 1.
@@
One of the interesting directions seems to be incorporating new types of explanations apart from visual ones (as it was done by Rabold et al. (2020) with symbolic rules). Moreover, it may be valuable to focus on studying the inner workings of the DL models and analyse them using some probe tasks (Adi et al., 2017). The motivation behind probe tasks is that if we can train a probe model on top of some internal representation (extracted from the DL model) to properly predict a property of the input data, it means that the property is encoded somewhere in the representation. Overall, we should not only focus on qualitative analyses but also try to enrich them with quantitative ones for a more holistic understanding of vision models’ decisions.
@@ 8 Conclusions
Recently, in the XAI field the shift of the paradigm from where to what was proposed meaning that instead of only highlighting some parts of input images, we should also be able to name them semantically. We propose to go further and focus on the notion of how the objects within the images are oriented towards each other (contextual information). We provide many examples when the ground truth labels depend on the spatial relationship between objects i.e. in autonomous cars, surveillance systems, healthcare, agriculture, and environment studies. In these cases, the existing XAI methods will fail to provide valid explanations of the DL models’ predictions. It may happen that despite having two different models’ predictions for two images, we would get semantically the same explanations.
We did a survey on papers targeting the topic of spatial context within DL models and XAI techniques. It turned out that there are only a few works in so-called spatial XAI compared to the number of DL models benefitting from spatial context. In this work, we would like to direct the attention of the XAI community to this largely unexplored area. We propose some potential research directions in the field of spatial XAI like a holistic benchmark that will evaluate models from the spatial context point of view and allow for quantitative assessment. We believe that by inspiring a community to address the topic of spatial XAI, it is possible to move forward research in XAI in general, as it would help in going out of the bubble of current development directions. Moreover, a better understanding of how the models ’see’ spatial context may lead to better DL models’ design in the future using a kind of reverse engineering method.
@@
Acknowledgements
The project was funded by POB Research Centre Cybersecurity and Data Science of Warsaw University of Technology within the Excellence Initiative Program - Research University (ID-UB). This research was carried out with the support of the Laboratory of Bioinformatics and Computational Genomics and the High Performance Computing Center of the Faculty of Mathematics and Information Science Warsaw University of Technology.

### v3

@@ 5 XAI Failures
We provide examples when popular XAI techniques fail to explain in a comprehensive way the predictions of the model when spatial context is crucial. We fine-tuned Resnet-50 and ViT base models in low data regime as it is done in work (Jia et al., 2022) using datasets from Visual Task Adaptation Benchmark (VTAB) (Zhai et al., 2019). The models were pretrained on Imagenet: Resnet-50 – in a supervised manner, ViT base – in a contrastive manner (Moco v3). We focused on the subset of VTAB called ‘structured’ where the labels depend on spatial context. In the experiments, we use two datasets: KITTI (Geiger et al., 2012) where images were collected using sensors in the car – the task is to predict the binned distance to the closest vehicle in the scene, dsprites (Matthey et al., 2017) where images of simple shapes undergo rotations and other shifts in the space – the task is to predict binned orientation. Hence, we analyze the only real-life dataset and one of the few synthetic datasets in VTAB.
Having fine-tuned models of a satisfactory performance (similar to the one claimed by Jia et al. (2022)), we applied 5 popular XAI techniques: GradientSHAP (Lundberg & Lee, 2017), Integrated gradients (Sundararajan et al., 2017), Occlusion (Zeiler & Fergus, 2014), Saliency (Simonyan et al., 2013), LIME (Ribeiro et al., 2016) to check if they manage to explain the model’s decisions correctly. In Figure 6, we provide images with the models’ explanations. The colors reflect the extent to which particular parts contribute to the final model’s decision (green – means positive attribution, whereas, red – negative). We can see that the existing popular XAI methods fail to explain the correct model decisions i.e. that the nearest vehicle is within the distance of 8 to 20 meters.
The explanations fail – they are inaccurate, vague and difficult to interpret (Figure 6). It can be seen that the nearest vehicle is highlighted in green (using the LIME method), however, it is not clear that the distance is crucial. One may conclude that the same explanations could also be seen in the object recognition tasks i.e. cars vs. planes. Therefore, it may be important to design spatial XAI techniques that would be tailored to different ML tasks or context types.
@@ 6 Spatial Context in XAI
The topic of spatial context was considered when training neural networks but it was of much smaller interest in XAI. We showed that there is still plenty to do as popular XAI methods fail. Note that without proper XAI tools, we would not be able to explain the predictions of CNNs and CapsuleNets when the same features are within input images but placed differently in the scene. The existing methods in the case of deformed face (Figure 5) will most probably highlight the same key elements of the image both in the case of CNNs and CapsuleNets even though the two DL models will return different predictions.
The investigation of whether the spatial context plays a role and to what extent in the decision-making of DL models is a niche that should be fulfilled. We call this area of research spatial XAI. There are only a few works in this domain. As a result of a survey, we distinguish four main approaches: intrinsically explainable models, methods introducing measures of spatial context, leveraging current XAI methods and the ones based on input-output relationships.
@@
The work (Bardool et al., 2019) encompasses two aspects as it proposes a new explainable-by-design architecture for detection and the metrics to evaluate the importance of context. The authors modify the well-known detector architecture by adding a contextual learner to extract contextual cues from the scenes. The contextual feature maps for each class give insights into the way the model benefits from context. The other contribution is the two scoring functions to measure the contextual relevance of detections in relation to other objects present in the image and the general scene. The first score is calculated as a summation of the elements of the feature map corresponding to a bounding box within the mask region, divided by the area of the bounding box’s mask. The second scoring method is analogous, but it is calculated based on contextual feature maps generated for each class instead of bounding boxes. The metrics are later used to compare different DL models. The authors challenged the proposed DL model on a set of hypotheses referring to spatial context. It was possible thanks to the fact that they had control over the generation of a synthetic dataset so that you could prepare samples violating the underlying spatial rules in the dataset and see how the model behaves in such scenarios.
@@
Sacha et al. (2023) analyze the spatial aspect within explainable-by-design prototypical models. They define the problem of ‘spatial explanation misalignment’ and introduce a metric to quantify this phenomenon. Moreover, they propose a benchmark and a method to mitigate this spatial misalignment in this particular kind of models. It was observed that sometimes the location of the explanation within the input image changes when the input image is modified in non-meaningful regions that should not influence the model prediction (distractors).
Tomaszewska et al. (2024) focuses on metrics to evaluate the preservation of spatial context in attention-based Multiple Instance Learning vision models that most often find application in the analysis of high-resolution images. The authors proposed three spatial context measures (, , ). The quantitative measurement is possible thanks to the incorporation of spatial regression models which are run on patch features to predict the respective attention scores assigned by the DL models. In the work, the histopathology use case is provided where the spatial context in the input is understood as spatial autocorrelation between patches of Whole Slide Images.
@@
One of the early XAI methods is LIME (Ribeiro et al., 2016) which operates on superpixels from input images assigning them the importance score explaining the model’s decision. In work, (Rabold et al., 2020), the authors propose to build upon the LIME method by incorporating logic rules obtained by the Inductive Logic Programming system, Aleph. The authors, therefore, propose to combine visual and symbolic methods to explain spatial relationships.
@@
In Deep Learning, often ablation studies are performed where the particular parts of the networks are switched off to check whether they are necessary and what their role is. Such an evaluation is often performed to compare the performance of the original network and the modified one. In XAI, a similar idea is behind permutation-based methods where parts of the input image are erased to see how it impacts predictions (i.e. LIME). In work, Tomaszewska et al. (2023), an analogous approach was used to evaluate the importance of neighborhood. The model used to predict whether a tissue patch contains a tumor or not was trained on images containing the region of interest and neighborhood. The ground truth label is provided only based on the region of interest. Later, inference was performed using images with a gradually decreased size of the neighborhood. It was analyzed how much the initial performance metrics were impacted by the limitation of contextual information. In the case of a so-defined neighborhood, a simple application of the existing XAI techniques could show whether models pay attention to the neighborhood or not.
Another input-output relationship approach accompanied by a new XAI method is proposed by Haldekar et al., 2017. The authors cover some parts of the input image (using a sliding window) and later compute the influence it has on a value of cross entropy compared to the situation when no modifications to the input image are made. As a result, the heatmaps showing key parts of the images are generated. The authors present some heatmaps visualizing the relationships (below, behind and above). However, the limitation of the provided heatmaps is the fact that without the authors’ captions, it would be difficult to decipher what type of relationships are highlighted. In the work, the authors also zero out different groups of neurons in MLP layers to check if they are responsible for detecting particular spatial relationships.
The analysis of spatial relation understanding is also performed in text-image generation models where it is assessed if the spatial relations specified in the prompt are preserved within generated images (Cho et al., 2022).
@@ 7 Needed Research Directions
In some of the mentioned papers on spatial context (Haldekar et al., 2017; Bardool et al., 2019; Swingler & Bath, 2020; Rabold et al., 2020), the claims were investigated on synthetically generated datasets in order to have better control over the experiment setup which facilitates the verification of the research hypotheses. These datasets were not made publicly available and, moreover, were not of sophisticated compositionality as there were only some simple objects in relative relationships put on the plain background. With the advent of diffusion models i.e. stable diffusion (Rombach et al., 2021), it is possible to generate more realistic images with the desired spatial relationships between objects (samples in Figure 3). Therefore, we postulate for new datasets reflecting scenarios where ground truth depends on spatial relationships.
Assessment of spatial reasoning skills in text-to-image models
PAINTSKILLS (Cho et al., 2022) is a compositional diagnostic evaluation dataset where images are generated using prompts containing information on spatial relationships (above, below, left, right) between objects. Note that these relationships are relatively simple compared to the ones we outlined in Section 2.2. The dataset is used as one of the scenarios verifying reasoning skills within the recently proposed benchmark, Holistic Evaluation of Text-to-Image Models (HEIM) (Lee et al., 2023). Within the reasoning scenario, it is checked whether models understand objects, counts, and spatial relations (compositionality) within images. Some task-specific benchmarks dedicated to the evaluation of the vision-language model’s ability to understand scene composition were proposed (Ma et al., 2022; Ray et al., 2023). It seems that the understanding of compositionality within text-to-image solutions is assessed only at a high level, however, in XAI the focus is on a deeper analysis of a model decision-making process. Here, we could conclude that the recent approaches are more about checking the accuracy of the model in outputting what is expected instead of tracking the model’s reasoning process.
@@
We believe that analogous or even more in-depth benchmarks should be proposed to verify the reasoning of purely vision models with a focus on assessing whether the spatial context is properly taken into account. To the best of our knowledge, there is no such benchmark even though there are many vision models that benefit from contextual information (Section 4). The analysis of some reasoning capabilities is performed in the case of models solving IQ tests but not in the real-world cases that we point out.
@@
A benchmark incorporating spatial relations understanding could serve as an important point in a checklist before the launch of DL models in critical domains. This could fulfill some of the requirements of ensuring the safety of models. For this benchmark to be successful, some new metrics to evaluate to what extent the models take into account spatial context should be proposed. So far models’ investigation mostly relies on visual XAI methods (heatmaps) which are not a sufficient way to explain models in cases when spatial context matters.
@@
One of the interesting directions seems to be incorporating new types of explanations on top of visual ones (as it was done by Rabold et al. (2020) with symbolic rules) or adding them as a separate element. Moreover, it may be valuable to focus on studying the inner workings of DL models and analyzing them using some probe tasks (Adi et al., 2017). The motivation behind probe tasks is that if we can train a probe model on top of some internal representation (extracted from the DL model) to correctly predict a property of the input data, it means that the property is encoded somewhere in the representation. Overall, we should not only focus on qualitative analyses but also try to enrich them with quantitative ones for a more holistic understanding of vision models’ decisions. In Appendix A, we discuss ideas on how to incorporate spatial XAI into a Data Science pipeline.
@@ 8 Conclusions
Recently, in the XAI field the shift of the paradigm from where to what was proposed, meaning that instead of only highlighting some parts of input images, we should also be able to name them semantically. We propose to go further and focus on the notion of how the objects within the images are oriented towards each other (contextual information). We provide many examples where the ground truth labels depend on the spatial relationship between objects i.e. in autonomous cars, surveillance systems, healthcare, agriculture, and environment studies. In these cases, the existing XAI methods will fail to provide valid explanations of the predictions of the DL models. It may happen that despite having two different models’ predictions for two images, we would get semantically the same explanations. We provide examples of failures of XAI methods (unclear explanations).
We conducted a survey on papers that focused on the topic of spatial context within DL models and XAI techniques. It turned out that there are only a few works in so-called spatial XAI compared to the number of DL models benefitting from spatial context. In this work, we would like to direct the attention of the XAI community to this largely unexplored area. We propose some potential research directions in the field of spatial XAI like a holistic benchmark that will evaluate models from the spatial context point of view and allow for quantitative assessment. We believe that by inspiring a community to address the topic of spatial XAI, it is possible to move forward research in XAI in general, as it would help in going out of the realm of current development directions. Moreover, a better understanding of how the models ’see’ spatial context may lead to better DL models’ design in the future.
Acknowledgements
We would like to thank the reviewers for the useful feedback that enriched the final version of the paper. This research was funded by the Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme. The work was carried out with the support of the Laboratory of Bioinformatics and Computational Genomics and the High Performance Computing Center of the Faculty of Mathematics and Information Science, Warsaw University of Technology.
@@ Acknowledgements
Appendix A Spatial XAI within Pipeline
The importance of spatial context within input data to DL models may not be black or white and some procedures may be needed when using spatial XAI in the future. There are many use cases where the spatial context is crucial. But there may be some exceptions. For instance, there can be an image of a room without any thermometer in it and we ask a model what the temperature in the room is. In this case, the context may not help in solving this task as simply visual information is not enough to solve this particular task.
We distinguish two possible scenarios of how to use the spatial XAI in the real-life Data Science pipeline but it may not be a finite list: (1) check in advance if the spatial XAI is applicable, (2) critically analyze the output of spatial XAI techniques.
In the first case, we assume that the process will be that the evaluators of the DL systems will know whether, in the particular task/ dataset, the spatial context is crucial. In some cases (i.e. ML for healthcare), the experts’ domain knowledge may be needed. This way it may be assessed if the use of spatial XAI techniques for the investigation of the model is justified. Such ‘check before use’ strategy is analogous to the case with the spatial regression models that can be applied instead of the Ordinary Least Square regression only if the observations and the residuals are not independent and form some clusters. This is formally checked from the perspective of spatial autocorrelation using the Global Moran’s I test. Maybe, a similar form of verification could be introduced to check if the use of spatial XAI method is justified in a particular case or not.
However, in some use cases, it may not be obvious if the spatial context is meaningful. Hence, the scenario could be that we apply the spatial XAI techniques to every use case and then evaluate the output – whether the spatial XAI methods show that indeed the spatial context plays a key role in the particular task. Such ‘post-hoc’ analysis can potentially bring some valuable insights into different phenomena where domain experts do not know a priori if the spatial context is important.